I'm having a hard time with this. Please help!
Using the quadratic equation x2 - 6x + 8 = 0, perform the following tasks:
a) Solve by factoring.
Answer:


Answer by checkley71(8403)

X^2-6X+8=0
(X-4)(X-2)=0
X-4=0
X=4 SOLUTION
X-2=0
X=2 SOLUTION